OpenStudy (anonymous):

Hey can someone help me with this question? What are the explicit equation and domain for an arithmetic sequence with a first term of 5 and a second term of 3?

OpenStudy (er.mohd.amir):

a1=5 and a2=3 then d=3-5=-2 a2=a1+d a3=a2+d and so on then AP is 5,3,1,-1,-3...............

OpenStudy (anonymous):

here are my answer choices an = 5 − 3(n − 1); all integers where n ≥ 1 an = 5 − 3(n − 1); all integers where n ≥ 0 an = 5 − 2(n − 1); all integers where n ≥ 0 an = 5 − 2(n − 1); all integers where n ≥ 1 Not sure how to find the right equation

OpenStudy (er.mohd.amir):

an=a1+(n-1)*d =5+(n-1)*(-2) =5-2(n-1),n>=1

OpenStudy (er.mohd.amir):

a1 is first term and a2 is second term and d is common difference =a2-a1 and formula an= a1+(n-1)*d put values.

OpenStudy (er.mohd.amir):

start n=1 since n=0 not give a1=5

OpenStudy (er.mohd.amir):

fine

OpenStudy (anonymous):

I see, thank you so much for your help